There’s a sign hanging from two ropes, the sign weighs 60kg. The angle in the top left is at 60 degrees and the angle in the top right is 30 degrees. What is tension in the wires
1
Expert's answer
2019-05-08T09:43:21-0400
"T_L \cos{60}=T_R \cos{30}"

"mg=T_L \sin{60}+T_R \sin{30}"

"mg=T_L \sin{60}+T_L \cos{60} \tan{30}=T_L (\sin{60}+\cos{60} \tan{30})"

"(60)(9.8)=T_L (\sin{60}+\cos{60} \tan{30})"

"T_L=509\ N""509 \cos{60}=T_R \cos{30}"

"T_R=294\ N"